OpenStudy (anonymous):

how do you write an equation from looking at a graph that already has the line drawn?

OpenStudy (anonymous):

use the slope intercept form (y = mx + b), where m represents the slope and b represents the y-intercept, look at the graph and find they y intercept, plug this value in for b, then find the slope and plug this value in for m

OpenStudy (anonymous):

ok oneBIG problem, no has ever show me how to read graphs so I don't have a clue what I am looking at.

OpenStudy (anonymous):

look at the graph where does it cross the y-axis ( the one that goes up and down)? that is what is substituted in for b then look at the line from the y-intercept, count how many units up and how many units over you have to go to get to the next point (this is the slope, plug it in for b), (example: if you count up 3 and right 5 the slope is 3/5)

OpenStudy (anonymous):

ok up 3 over to the right 7 =3/7 right? then how do i figure for y=?

OpenStudy (anonymous):

y=mx + b, where m = slope, and b = y-intercept

OpenStudy (anonymous):

ok m = 3/5 and b=3???

OpenStudy (anonymous):

so y=3/5x + 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cool I got one right. thank you very much

OpenStudy (anonymous):

congratulations

OpenStudy (anonymous):

i had another one to do and the way you explained it I was able to do that one rather quickly

OpenStudy (anonymous):

now if u can help me do the equation and then draw the line on a blank graph y=-2x+5

OpenStudy (anonymous):

m = -2, and b = 5 so go up 5 units on the y-axis and put a point, then count from the point (up 2, left 1) or (down 2, right 1) to find more points, then connect them to graph the line

OpenStudy (anonymous):

ok how do u know if u go up 2 or down 2?

OpenStudy (anonymous):

if the slope is negative the line will tilt up on the left and down on the right, so (up 2, and left 1) will the same as (down 2 and right 1)

OpenStudy (anonymous):

ok i see. thanks. now the next one is different. x+5y=10

OpenStudy (anonymous):

you need to solve for y, so move the x to the right (5y = -x + 10), then divide all the terms by 5 (y = -1/5x + 2), so the y-intercept is 2, and the slope is -1/5 (up 1, left 5 or down 1, right 5)

OpenStudy (anonymous):

ok thanks